# MnLargeSymbols’164 MnLargeSymbols’171

### v2

@@ Appendix A Nonequilibrium behavior
In Fig. 7a) an hysteresis cycle is shown for MC simulations of the IMB/TMB case for one ER network of size $N=64$ and average degree $\bar{k}=8$. The cycle begins at $T=0$ starting from an uniform initial condition, $x_{i}=1/N$. As can be seen in the figure, a loop between the temperatures $T=0$ and $T_{c}$, where the curve during the heathing phase (right triangles) goes mostly below the curve of the cooling phase (left triangles). This results differs from the observed in Fig. 2, indicating that the system dynamics is not at thermodynamic equilibrium and, therefore, the initial condition plays an important role. The area within the loop varies significantly with the network sample, but decreases with NN possibly due to the self-averaging of the network heterogeneities (compare Figs. 7 a) and b)). Interestingly, the curves exhibit a region where the average share of wealth $\left\langle x_{1}\right\rangle_{a}$ of the richest agent grows with $T$. This counterintuitive behavior is shared by the other rich agents (see Fig.7b)). Moreover, as shown in Fig. 8, the opposite tendency is observed for the number of rich agents. These results, together with the absence of the plateau in QMF approximations, suggest that local wealth condensation is a necessary but not sufficient condition for the discrepancy to occur.
As depicted in Fig. 9, a similar behavior is also found for IMA/TMA, but with a significantly smaller hysteresis loop. More specifically, the area within the loop decreases with NN like in the TMB cases (see Fig. 9a)), but grows with the average degree $\bar{k}$ (see Fig. 9b)) where, to facilitate the visual comparison for the different sizes NN, the curves are adjusted by the scaling of Eq. 15.
@@
Figure 9: Scaling of hysteresis effects with connectivity and system size. a) Hysteresis cycles of $N^{1/4}x_{1}$ for different system sizes NN, with fixed mean connectivity at $\bar{k}/N\approx 0.25$. b) Hysteresis cycles of $x_{1}$ for $N=64$ and varying average degree $\bar{k}$. Data are taken from the same simulations as in Fig. 8b).

### v3

@@ Appendix A Nonequilibrium behavior
In Fig. 7a) an hysteresis cycle is shown for MC simulations of the IMB/TMB case for one ER network of size $N=64$ and average degree $\bar{k}=8$. The cycle begins at $T=0$ starting from an uniform initial condition, $x_{i}=1/N$. As can be seen in the figure, a loop between the temperatures $T=0$ and $T_{c}$, where the curve during the heathing phase (right triangles) goes mostly below the curve of the cooling phase (left triangles). This results differs from the observed in Fig. 2, indicating that the system dynamics is not at thermodynamic equilibrium and, therefore, the initial condition plays an important role. The area within the loop varies significantly with the network sample, but decreases with $N$ possibly due to the self-averaging of the network heterogeneities (compare Figs. 7 a) and b)). Interestingly, the curves exhibit a region where the average share of wealth $\left\langle x_{1}\right\rangle_{a}$ of the richest agent grows with $T$. This counterintuitive behavior is shared by the other rich agents (see Fig.7b)). Moreover, as shown in Fig. 8, the opposite tendency is observed for the number of rich agents. These results, together with the absence of the plateau in QMF approximations, suggest that local wealth condensation is a necessary but not sufficient condition for the discrepancy to occur.
As depicted in Fig. 9, a similar behavior is also found for IMA/TMA, but with a significantly smaller hysteresis loop. More specifically, the area within the loop decreases with $N$ like in the TMB cases (see Fig. 9a)), but grows with the average degree $\bar{k}$ (see Fig. 9b)) where, to facilitate the visual comparison for the different sizes $N$, the curves are adjusted by the scaling of Eq. 15.
@@
Figure 9: Scaling of hysteresis effects with connectivity and system size. a) Hysteresis cycles of $N^{1/4}x_{1}$ for different system sizes $N$, with fixed mean connectivity at $\bar{k}/N\approx 0.25$. b) Hysteresis cycles of $x_{1}$ for $N=64$ and varying average degree $\bar{k}$. Data are taken from the same simulations as in Fig. 8b).